# How Far Are We from Optimal Reasoning Efficiency?

### v1

@@ 1 Introduction
Large Reasoning Models (LRMs) have recently emerged as a powerful class of models capable of solving complex tasks that require advanced reasoning. Frontier LRMs such as OpenAI o1 [31] and DeepSeek R1 [9] have obtained superior performance across a wide range of tasks, including mathematical reasoning and competitive programming. A major factor behind this success is their ability to perform deep, multi-step reasoning through extended Chain-of-Thought (CoT) processes. These reasoning traces often include sophisticated operations such as reflection, verification, and exploration, within a single inference pass.
However, the powerful capability of long CoT reasoning comes at a cost. LRMs frequently generate overly verbose and redundant reasoning traces, a phenomenon referred to as the overthinking problem [51, 40]. Recent studies [4, 40] have shown that even simple questions like “2 + 3 = ?” can result in outputs spanning up to 900 tokens. This redundancy brings a significant cost in inference time and limits practical deployment. Several fine-tuning approaches have been proposed to improve reasoning efficiency, with representative examples being length reduction with RL [43, 24, 1, 2, 52, 38, 34, 50, 36, 12], and hybrid reasoning approaches that train LRMs to adaptively select thinking or no-thinking modes [15, 23, 58, 19]. However, comparing these methods remains difficult due to inconsistent experiment setups, including varying models, evaluation benchmarks, and mixed performance metrics. It is still unclear how close current approaches are to the optimal trade-off between length and accuracy.
@@ 2 Related Works
Prior studies have shown that LRMs often suffer from redundant reasoning. Even for very simple questions, Frontier LRMs often generate lengthy responses spanning thousands of tokens [5, 40]. This redundancy in the reasoning process brings significant overheads in the inference costs. Several works are then proposed to make LRM reasoning more concise. [43, 24, 1, 2, 52, 38, 34, 50, 36, 12, 33] investigate RL training with length reward designs, mostly focusing on reducing the reasoning lengths. A line of works apply SFT to fine-tune LRMs on datasets with variable-length reasoning traces to elicit concise reasoning [54, 45, 10] or adjustable length control [16, 47, 27, 21, 56]. Some recent works investigate training LRMs to adaptively select whether to perform an explicit thinking process or directly produce the final solution according to the input query [58, 44, 15]. Test-time techniques are also applied for enhancing the reasoning efficiency through reward model guided decoding [41, 20], uncertainty-based dynamic reasoning [7, 8], and confidence-based approaches [42, 14]. In this work, we study the optimal reasoning efficiency for LRMs and focus on training-based approaches for enhancing LRM reasoning efficiency. Our method aims at enhancing the accuracy of the LRM under diverse token budgets without explicitly incentivizing shorter responses.
@@
Reinforcement Learning is the central technique for eliciting and enhancing the reasoning capability of LRMs. Frontier LRMs, including OpenAI o1 [31] and DeepSeek R1 [9], have shown that applying "zero RL" on a base LLM could effectively elicit the ability to utilize long CoTs for complex reasoning. A series of works have emerged with the focus on improving the training efficiency of RL for LRMs from the perspectives of data [26, 35, 11, 18, 46], algorithms [9, 11, 26, 55, 57], and training framework [39, 25, 35]. A number of works successfully apply zero RL training on a wide range of reasoning-heavy domains, including multi-modality [37, 59], medical [53, 3], and financial [22]. Recent works also explore efficiency enhancement by encouraging concise reasoning with RL [43, 24, 1, 2, 52, 38, 34, 50, 36, 12]. In this work, we focus on enhancing the reasoning efficiency with RL.
@@ 4.1 Defining Optimality in Token-Bounded Reasoning
To assess optimal reasoning efficiency, we must evaluate the performance of an LRM $\pi_{\theta}$ under a fixed token budget $L$. Simply truncating an output after $L$ tokens, however, may lead to incomplete responses with no answers. To address this, we define a fallback mechanism following prior works [29, 7]. If the reasoning trace $y\sim\pi_{\theta}(\cdot|x)$ exceeds $L$ tokens, the model is prompted to produce a final answer $a$ directly from the truncated trace $y_{:L}$,
@@ 4.2 Empirical Estimation of Reasoning Frontiers
  • Online RL with Token Budgets: We conduct online RL training with different token budgets ranging from 512 to 32k. Fine-tuning an LRM with a token budget effectively enforces the LRM to reason with limited tokens [48, 12].
  • Online RL with Length Rewards. We test various reward designs that promote concise yet accurate reasoning, including length-harmonizing rewards [24] and length-group normalized rewards [2].
  • Preference Learning. We apply SimPO [28] with preference datasets constructed via methods such as TOPS [51] and DAST [38]. For example, we contrast short correct responses with longer ones to promote conciseness [30].
@@
We conduct our study using two classes of base LRMs, DeepSeek-R1-Distill-Qwen-1.5B/7B [9] and Qwen3-4B/8B [49]. Specially, for DeepSeek-R1-Distill-Qwen-1.5B/7B, we start from RL-fine-tuned versions and further fine-tune them using the strategies above. The evaluation is carried out on a set of challenging mathematical reasoning benchmarks. More training and evaluation details can be found in Sec. 6.1 and Appendix. C.
@@
As a single metric, REG unifies both the accuracy and length aspects. At the same time, REG quantitatively assesses the gap of reasoning efficiency between an LRM and efficiency frontiers. Most of the existing evaluation metrics only focus on one of the accuracy and length aspects. For example, Pu et al. [32] proposes to evaluate reasoning redundancy by computing the length difference between the longest and the shortest response, without considering correctness of the generated responses. We also note that Chen et al. [4] introduces an outcome efficiency metric that measures the ratio between the position of first reaching the correct answer and the the overall length for correct responses. Although this outcome efficiency metric takes both accuracy and length into consideration, it has the risk of being hacked if the LRM only produces an answer at the end of the response.
@@ 6.1 Experimental Setup
We use DeepSeek-R1-Distill-Qwen-1.5B&7B and Qwen3-4B&8B as the base LRMs. For training, we adopt a mixture of training data consisting of 135k problems sourced from DeepScaleR [26] and AReaL [35] For evaluation, we use several challenging mathematical benchmarks: AMC 2023, AIME 2024 & 2025, and Minerva Math [17]. We report the average accuracy of 32 responses generated with temperature $T=0.6$ and top_p$=0.95$ with maximum length $L_{\max}=32K$. The main results are averaged over three benchmarks. When evaluating REG, we use a smaller length $L_{\text{max}}=16K$.
@@
  • Online RL with Length-based Rewards: We consider online RL with length-based rewards, including length-harmonizing rewards [24], and length group normalized rewards [2]. We make comparison with Meta Reinforcement Fine-tuning (MRT) [34], which minimizes regret for individual steps to enhance reasoning efficiency. Additionally, we also adopt online RL with hard token budgets as a baseline, which is also adopted in recent works [12, 48]. We note that a recent work BRPO [33] optimizes rewards under linearly spaced token budgets, which is similar to a variant of REO-RL studied in the ablation study (Sec. 6.4). [^3] A more detailed discussion between BRPO and REO-RL is provided in Appendix. A
  • Hybrid Reasoning. We use HGPO [15] as a representative hybrid reasoning baseline that trains the LRM to adaptively select between the thinking and no-thinking modes. There are also recent works sharing the similar idea of encouraging no-thinking mode when no-thinking mode achieves competitive accuracy [6, 58, 44].
  • Supervised Fine-Tuning: For each problem in the training data, we generate multiple responses and perform SFT on the shortest correct one. Two strategies are used for data generation. The first strategy is direct generation with the problems as inputs [30]. The second strategy follows TOPS to prompt the LRM with different levels of reasoning efforts [51].
  • Preference Learning: We apply SimPO [28] on various preference datasets. In $SimPO_{shortest}$, we use the shortest correct response and the longest response as the preference pairs. We also follow TOPS [51] and DAST [38] to construct preference datasets.
We have also tried RL-based length control methods such as  [1, 48] but find these methods only achieve successful length control under low token budgets and yield similar length-accuracy trade-off as RL w. Token Budgets. More details about the baselines and our other investigations could be found in Appendix. D.
@@
For REO-RL and all baselines, we implement all methods based on the AReaL framework [35]. For DeepSeek-R1-Distill-Qwen-1.5B/7B, we fine-tune the corresponding RL-trained versions. Specifically, we use AReaL-Boba-RL-1.5B [35] and SkyWork-OR1-Math-7B [11] as the starting points for further fine-tuning in 1.5B and 7B experiments, respectively. For Qwen3 models, we directly apply RL training on the base LRMs. For REO-RL (Exp) and REO-RL (Oracle), we use $N=5$ token budgets since $N=5$ already obtains sufficiently accurate approximations as illustrated in Fig. 3. For more training details, please refer to Appendix. C.
@@ 6.2 Main Results
RL w. Token Budget=1K 17.3↓24.9 1487.4 $\uparrow{170.4\%}$ 13.5↓14.7 1282.8 $\uparrow{140.7\%}$ 65.6↓16.2 $1$.3 $\uparrow{43.1\%}$ 32.0↓3.2 1108.8 $\downarrow{24.1\%}$ 32.1↓14.8 1247.6 $\uparrow{90.3\%}$
@@
REO-RL (Oracle) (ours) 63.9↓2.4 9348.6 $\downarrow{50.6\%}$ 49.0↓4.0 9189.6 $\downarrow{66.5\%}$ 94.7↑0.8 $4$.1 $\downarrow{57.9\%}$ 45.5↑0.9 4132.9 $\downarrow{57.8\%}$ 63.3↓1.2 6778.8 $\downarrow{58.0\%}$
@@ 6.4 Ablation Study of REO-RL
REO-RL (Oracle) 63.9↓2.4 9348.6 $\downarrow{50.6\%}$ 49.0↓4.0 9189.6 $\downarrow{66.5\%}$ 94.7↑0.8 $4$.1 $\downarrow{57.9\%}$ 45.5↑0.9 4132.9 $\downarrow{57.8\%}$ 63.3↓1.2 6778.8 $\downarrow{58.0\%}$
@@ Appendix C Implementation Details
For training data, we integrate data from DeepScaleR [26] and AReaL [35]. For models with at least 4B sizes, we use the training data of AReaL-Boba-RL-7B [35]. For 1.5B, we adopt the mixture of training data from DeepScaleR [26] and AReaL [35] and remove duplicated problems.
We implement the training algorithm with the AReaL framework [35], which supports SGLang [60] for rollout generation. Below we detail implementation details of each training algorithm.
@@
We use PPO as the default online RL algorithm. Following standard practices in RL training for LLM reasoning [55, 13], we do not utilize value model and KL regularization. The default training setting and hyperparameters for PPO training are listed in Tab. 4.
For REO-RL and baseline methods, we do not carry out RL training directly from the base LRM since it would lead to prolonged training and slower convergence. Instead, we perform further fine-tuning on the RL trained versions for both 1.5B and 7B settings. Specifically, we adopt AReaL-Boba-RL-1.5B [35] and Skywork-OR1-Math-7B [11] as the starting points for further fine-tuning. Following the cluster configuration in Tab. 4, each experiment could finish within 4K GPU hours.
@@
We implement SimPO [28] in the AReaL framework [35]. The default training configurations and hyperparameters for SimPO are listed in Tab. 5.
@@ Appendix D Baselines
In the “RL with Length Group Normalized Rewards” baseline [2], for each question $x$ and the corresponding set of sampled responses $y^{1},\cdots,y^{m}$, the reward of response $y^{i}$ is computed as,
@@
In the “RL with Length-Harmonizing Rewards” baseline [24], for each question $x$ and the corresponding set of sampled responses $y^{1},\cdots,y^{m}$, the reward of response $y^{i}$ is computed as,
@@
In the MRT baseline [34], different from the original paper that only implements single-step optimization with offline collected response prefixes, we implement the online RL training version with dense rewards for MRT. For each question $x$ and the corresponding set of sampled responses , we partition each response into several steps $y=(y_{1},\cdots,y_{s})$. In each training step, the model is updated by computing the policy gradient for the following objective,
@@
We use “\$\backslash n$” and “\$\backslash n$\$\backslash n$</think>” as the tokens to start thinking and no-thinking. We first run SFT over model-generated responses in the thinking and no-thinking modes to ensure the model initially has a non-zero probability for no-thinking modes. Then we apply HGPO training with a margin of $0.2$ following  [15].
@@
In $\text{SFT}_{\text{Shortest}}$, we generate $16$ outputs for each question in the training dataset. Then we select the correct response with the shortest length for each question to construct the SFT dataset. In $\text{SFT}_{\text{TOPS}}$, we follow [51] to prompt the LRM to generate responses with three different types of reasoning efforts. We strictly follow the prompts used in [51]. For each type of reasoning effort, we generate $16$ responses. To construct the SFT dataset, the shortest correct response among all $48$ responses are gathered.
@@
We have also experimented with Z1 [56] that constructs code-based reasoning traces, and L1 [1] that fine-tunes the LRMs to follow token budget instruction with RL. We find Z1 having poor performance on mathematical reasoning tasks, demonstrating significantly lower accuracy to that of the base LRMs as illustrated in Tab. 7. For L1, we find L1 mainly works under tight token budgets, i.e. less than $4K$. When we extend the training approach of L1 to a larger context length, i.e. 24K for 1.5B models, we find it hard to make the LRM learning to follow strict token budget instructions through RL. Consequently, the resulted models, L1-Exact-24K-1.5B and L1-Max-24K-1.5B could not follow the token budget instruction, as shown in Tab, 8.
@@
Table 7: Result of Z1 on DeepSeek-R1-Distill-Qwen-1.5B. [56]

### v2

@@ 1 Introduction
Large Reasoning Models (LRMs) have recently emerged as a powerful class of models capable of solving complex tasks that require advanced reasoning. Frontier LRMs such as OpenAI o1 (OpenAI, ) and DeepSeek R1 (Guo et al., 2025) have obtained superior performance across a wide range of tasks, including mathematical reasoning and competitive programming. A major factor behind this success is their ability to perform deep, multi-step reasoning through extended Chain-of-Thought (CoT) processes. These reasoning traces often include sophisticated operations such as reflection, verification, and exploration, within a single inference pass.
However, the powerful capability of long CoT reasoning comes at a cost. LRMs frequently generate overly verbose and redundant reasoning traces, a phenomenon referred to as the overthinking problem (Yang et al., 2025c; Sui et al., 2025). Recent studies (Chen et al., 2024b; Sui et al., 2025) have shown that even simple questions like “2 + 3 = ?” can result in outputs spanning up to 900 tokens. This redundancy brings a significant cost in inference time and limits practical deployment. Several fine-tuning approaches have been proposed to improve reasoning efficiency, with representative examples being length reduction with RL (Team et al., 2025; Luo et al., 2025a; Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Yeo et al., 2025; Shen et al., 2025b; Qu et al., 2025; Yang et al., 2025b; She et al., 2025; Hou et al., 2025), and hybrid reasoning approaches that train LRMs to adaptively select thinking or no-thinking modes (Jiang et al., 2025; Lou et al., 2025; Zhang et al., 2025a; Liang et al., 2025). However, comparing these methods remains difficult due to inconsistent experiment setups, including varying models, evaluation benchmarks, and mixed performance metrics. It is still unclear how close current approaches are to the optimal trade-off between length and accuracy.
@@ 2 Related Works
Prior studies have shown that LRMs often suffer from redundant reasoning. Even for very simple questions, Frontier LRMs often generate lengthy responses spanning thousands of tokens (Chen et al., 2025; Sui et al., 2025). This redundancy in the reasoning process brings significant overheads in the inference costs. Several works are then proposed to make LRM reasoning more concise. Team et al. (2025); Luo et al. (2025a); Aggarwal and Welleck (2025); Arora and Zanette (2025); Yeo et al. (2025); Shen et al. (2025b); Qu et al. (2025); Yang et al. (2025b); She et al. (2025); Hou et al. (2025); Qi et al. (2025) investigate RL training with length reward designs, mostly focusing on reducing the reasoning lengths. A line of works apply SFT to fine-tune LRMs on datasets with variable-length reasoning traces to elicit concise reasoning (Yu et al., 2024; Wang et al., 2023; Han et al., 2024) or adjustable length control (Kang et al., 2024; Xia et al., 2025; Ma et al., 2025; Liu et al., 2024; Yu et al., 2025c). Some recent works investigate training LRMs to adaptively select whether to perform an explicit thinking process or directly produce the final solution according to the input query (Zhang et al., 2025a; Tu et al., 2025; Jiang et al., 2025). Test-time techniques are also applied for enhancing the reasoning efficiency through reward model guided decoding (Sun et al., 2024; Liao et al., 2025), uncertainty-based dynamic reasoning (Fu et al., 2024, 2025), and confidence-based approaches (Taubenfeld et al., 2025; Huang et al., 2025). In this work, we study the optimal reasoning efficiency for LRMs and focus on training-based approaches for enhancing LRM reasoning efficiency. Our method aims at enhancing the accuracy of the LRM under diverse token budgets without explicitly incentivizing shorter responses.
@@
Reinforcement Learning is the central technique for eliciting and enhancing the reasoning capability of LRMs. Frontier LRMs, including OpenAI o1 (OpenAI, ) and DeepSeek R1 (Guo et al., 2025), have shown that applying "zero RL" on a base LLM could effectively elicit the ability to utilize long CoTs for complex reasoning. A series of works have emerged with the focus on improving the training efficiency of RL for LRMs from the perspectives of data (Luo et al., 2025c; RL Lab, 2025; He et al., 2025; Li et al., 2025; Wang et al., 2025), algorithms (Guo et al., 2025; He et al., 2025; Luo et al., 2025c; Yu et al., 2025b; Yue et al., 2025), and training framework (Sheng et al., 2024; Luo et al., 2025b; RL Lab, 2025). A number of works successfully apply zero RL training on a wide range of reasoning-heavy domains, including multi-modality (Shen et al., 2025a; Zhang et al., 2025b), medical (Yu et al., 2025a; Chen et al., 2024a), and financial (Liu et al., 2025). Recent works also explore efficiency enhancement by encouraging concise reasoning with RL (Team et al., 2025; Luo et al., 2025a; Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Yeo et al., 2025; Shen et al., 2025b; Qu et al., 2025; Yang et al., 2025b; She et al., 2025; Hou et al., 2025). In this work, we focus on enhancing the reasoning efficiency with RL.
@@ 4.1 Defining Optimality in Token-Bounded Reasoning
To assess optimal reasoning efficiency, we must evaluate the performance of an LRM $\pi_{\theta}$ under a fixed token budget $L$. Simply truncating an output after $L$ tokens, however, may lead to incomplete responses with no answers. To address this, we define a fallback mechanism following prior works (Muennighoff et al., 2025; Fu et al., 2024). If the reasoning trace $y\sim\pi_{\theta}(\cdot|x)$ exceeds $L$ tokens, the model is prompted to produce a final answer $a$ directly from the truncated trace $y_{:L}$,
@@ 4.2 Empirical Estimation of Reasoning Frontiers
  • Online RL with Token Budgets: We conduct online RL training with different token budgets ranging from 512 to 32k. Fine-tuning an LRM with a token budget effectively enforces the LRM to reason with limited tokens (Xu et al., 2025; Hou et al., 2025).
  • Online RL with Length Rewards. We test various reward designs that promote concise yet accurate reasoning, including length-harmonizing rewards (Luo et al., 2025a) and length-group normalized rewards (Arora and Zanette, 2025).
  • Preference Learning. We apply SimPO (Meng et al., 2024) with preference datasets constructed via methods such as TOPS (Yang et al., 2025c) and DAST (Shen et al., 2025b). For example, we contrast short correct responses with longer ones to promote conciseness (Munkhbat et al., 2025).
@@
We conduct our study using two classes of base LRMs, DeepSeek-R1-Distill-Qwen-1.5B/7B (Guo et al., 2025) and Qwen3-4B/8B (Yang et al., 2025a). Specially, for DeepSeek-R1-Distill-Qwen-1.5B/7B, we start from RL-fine-tuned versions and further fine-tune them using the strategies above. The evaluation is carried out on a set of challenging mathematical reasoning benchmarks. More training and evaluation details can be found in Sec. 6.1 and Appendix. C.
@@
As a single metric, REG unifies both the accuracy and length aspects. At the same time, REG quantitatively assesses the gap of reasoning efficiency between an LRM and efficiency frontiers. Most of the existing evaluation metrics only focus on one of the accuracy and length aspects. For example, Pu et al. (2025) proposes to evaluate reasoning redundancy by computing the length difference between the longest and the shortest response, without considering correctness of the generated responses. We also note that Chen et al. (2024b) introduces an outcome efficiency metric that measures the ratio between the position of first reaching the correct answer and the the overall length for correct responses. Although this outcome efficiency metric takes both accuracy and length into consideration, it has the risk of being hacked if the LRM only produces an answer at the end of the response.
@@ 6.1 Experimental Setup
We use DeepSeek-R1-Distill-Qwen-1.5B&7B and Qwen3-4B&8B as the base LRMs. For training, we adopt a mixture of training data consisting of 135k problems sourced from DeepScaleR Luo et al. (2025c) and AReaL RL Lab (2025) For evaluation, we use several challenging mathematical benchmarks: AMC 2023, AIME 2024 & 2025, and Minerva Math (Lewkowycz et al., 2022). We report the average accuracy of 32 responses generated with temperature $T=0.6$ and top_p=0.95=0.95 with maximum length $L_{\max}=32K$. The main results are averaged over three benchmarks. When evaluating REG, we use a smaller length $L_{\text{max}}=16K$.
@@
  • Online RL with Length-based Rewards: We consider online RL with length-based rewards, including length-harmonizing rewards (Luo et al., 2025a), and length group normalized rewards (Arora and Zanette, 2025). We make comparison with Meta Reinforcement Fine-tuning (MRT) (Qu et al., 2025), which minimizes regret for individual steps to enhance reasoning efficiency. Additionally, we also adopt online RL with hard token budgets as a baseline, which is also adopted in recent works (Hou et al., 2025; Xu et al., 2025). We note that a recent work BRPO (Qi et al., 2025) optimizes rewards under linearly spaced token budgets, which is similar to a variant of REO-RL studied in the ablation study (Sec. 6.4). [^3] A more detailed discussion between BRPO and REO-RL is provided in Appendix. A
  • Hybrid Reasoning. We use HGPO (Jiang et al., 2025) as a representative hybrid reasoning baseline that trains the LRM to adaptively select between the thinking and no-thinking modes. There are also recent works sharing the similar idea of encouraging no-thinking mode when no-thinking mode achieves competitive accuracy (Fang et al., 2025; Zhang et al., 2025a; Tu et al., 2025).
  • Supervised Fine-Tuning: For each problem in the training data, we generate multiple responses and perform SFT on the shortest correct one. Two strategies are used for data generation. The first strategy is direct generation with the problems as inputs (Munkhbat et al., 2025). The second strategy follows TOPS to prompt the LRM with different levels of reasoning efforts (Yang et al., 2025c).
  • Preference Learning: We apply SimPO (Meng et al., 2024) on various preference datasets. In $SimPO_{shortest}$, we use the shortest correct response and the longest response as the preference pairs. We also follow TOPS (Yang et al., 2025c) and DAST (Shen et al., 2025b) to construct preference datasets.
We have also tried RL-based length control methods such as  (Aggarwal and Welleck, 2025; Xu et al., 2025) but find these methods only achieve successful length control under low token budgets and yield similar length-accuracy trade-off as RL w. Token Budgets. More details about the baselines and our other investigations could be found in Appendix. D.
@@
For REO-RL and all baselines, we implement all methods based on the AReaL framework (RL Lab, 2025). For DeepSeek-R1-Distill-Qwen-1.5B/7B, we fine-tune the corresponding RL-trained versions. Specifically, we use AReaL-Boba-RL-1.5B (RL Lab, 2025) and SkyWork-OR1-Math-7B (He et al., 2025) as the starting points for further fine-tuning in 1.5B and 7B experiments, respectively. For Qwen3 models, we directly apply RL training on the base LRMs. For REO-RL (Exp) and REO-RL (Oracle), we use $N=5$ token budgets since $N=5$ already obtains sufficiently accurate approximations as illustrated in Fig. 3. For more training details, please refer to Appendix. C.
@@ 6.2 Main Results
RL w. Token Budget=1K 17.3↓24.9 1487.4 $\uparrow{170.4\%}$ 13.5↓14.7 1282.8 $\uparrow{140.7\%}$ 65.6↓16.2 1111.3 $\uparrow{43.1\%}$ 32.0↓3.2 1108.8 $\downarrow{24.1\%}$ 32.1↓14.8 1247.6 $\uparrow{90.3\%}$
@@
REO-RL (Oracle) (ours) 63.9↓2.4 9348.6 $\downarrow{50.6\%}$ 49.0↓4.0 9189.6 $\downarrow{66.5\%}$ 94.7↑0.8 4444.1 $\downarrow{57.9\%}$ 45.5↑0.9 4132.9 $\downarrow{57.8\%}$ 63.3↓1.2 6778.8 $\downarrow{58.0\%}$
@@ 6.4 Ablation Study of REO-RL
REO-RL (Oracle) 63.9↓2.4 9348.6 $\downarrow{50.6\%}$ 49.0↓4.0 9189.6 $\downarrow{66.5\%}$ 94.7↑0.8 4444.1 $\downarrow{57.9\%}$ 45.5↑0.9 4132.9 $\downarrow{57.8\%}$ 63.3↓1.2 6778.8 $\downarrow{58.0\%}$
@@ Appendix C Implementation Details
For training data, we integrate data from DeepScaleR Luo et al. (2025c) and AReaL RL Lab (2025). For models with at least 4B sizes, we use the training data of AReaL-Boba-RL-7B (RL Lab, 2025). For 1.5B, we adopt the mixture of training data from DeepScaleR Luo et al. (2025c) and AReaL RL Lab (2025) and remove duplicated problems.
We implement the training algorithm with the AReaL framework (RL Lab, 2025), which supports SGLang (Zheng et al., 2024) for rollout generation. Below we detail implementation details of each training algorithm.
@@
We use PPO as the default online RL algorithm. Following standard practices in RL training for LLM reasoning (Yu et al., 2025b; Hu et al., 2025), we do not utilize value model and KL regularization. The default training setting and hyperparameters for PPO training are listed in Tab. 4.
For REO-RL and baseline methods, we do not carry out RL training directly from the base LRM since it would lead to prolonged training and slower convergence. Instead, we perform further fine-tuning on the RL trained versions for both 1.5B and 7B settings. Specifically, we adopt AReaL-Boba-RL-1.5B (RL Lab, 2025) and Skywork-OR1-Math-7B (He et al., 2025) as the starting points for further fine-tuning. Following the cluster configuration in Tab. 4, each experiment could finish within 4K GPU hours.
@@
We implement SimPO (Meng et al., 2024) in the AReaL framework (RL Lab, 2025). The default training configurations and hyperparameters for SimPO are listed in Tab. 5.
@@ Appendix D Baselines
In the “RL with Length Group Normalized Rewards” baseline (Arora and Zanette, 2025), for each question $x$ and the corresponding set of sampled responses $y^{1},\cdots,y^{m}$, the reward of response $y^{i}$ is computed as,
@@
In the “RL with Length-Harmonizing Rewards” baseline (Luo et al., 2025a), for each question $x$ and the corresponding set of sampled responses $y^{1},\cdots,y^{m}$, the reward of response $y^{i}$ is computed as,
@@
In the MRT baseline (Qu et al., 2025), different from the original paper that only implements single-step optimization with offline collected response prefixes, we implement the online RL training version with dense rewards for MRT. For each question $x$ and the corresponding set of sampled responses , we partition each response into several steps $y=(y_{1},\cdots,y_{s})$. In each training step, the model is updated by computing the policy gradient for the following objective,
@@
[…]
@@
In $\text{SFT}_{\text{Shortest}}$, we generate $16$ outputs for each question in the training dataset. Then we select the correct response with the shortest length for each question to construct the SFT dataset. In $\text{SFT}_{\text{TOPS}}$, we follow (Yang et al., 2025c) to prompt the LRM to generate responses with three different types of reasoning efforts. We strictly follow the prompts used in (Yang et al., 2025c). For each type of reasoning effort, we generate $16$ responses. To construct the SFT dataset, the shortest correct response among all $48$ responses are gathered.
@@
We have also experimented with Z1 (Yu et al., 2025c) that constructs code-based reasoning traces, and L1 (Aggarwal and Welleck, 2025) that fine-tunes the LRMs to follow token budget instruction with RL. We find Z1 having poor performance on mathematical reasoning tasks, demonstrating significantly lower accuracy to that of the base LRMs as illustrated in Tab. 7. For L1, we find L1 mainly works under tight token budgets, i.e. less than $4K$. When we extend the training approach of L1 to a larger context length, i.e. 24K for 1.5B models, we find it hard to make the LRM learning to follow strict token budget instructions through RL. Consequently, the resulted models, L1-Exact-24K-1.5B and L1-Max-24K-1.5B could not follow the token budget instruction, as shown in Tab, 8.
@@
Table 7: Result of Z1 on DeepSeek-R1-Distill-Qwen-1.5B. (Yu et al., 2025c)